OpenStudy (anonymous):

.Jess uses her office scanner to scan pages at the rate of $0.09 per page. She decides to rent a scanner for $70 a year.The cost of scanning used rented the rented scanner is $0.02 per page. Part A: write an inequality that can be used to calculate the number of pages that Jess should scan in a year so that the amount she pays for the rented scanner is less than the office scanner. Define the variable used

OpenStudy (austinl):

We have two different copiers. Copier 1 and copier 2. Copier 1 costs 0.09 per page. Coper 2 costs 0.02 per page, plus a flat fee of 70. Let's use x as our variable, where x=pages printed. We want to find the number of pages that she needs to scan in a year that would make copier 2 more cost effective. So, as such we have, \(0.02x+70<0.09x\) We would need to solve for x, do you think you could do that for me?

OpenStudy (anonymous):

oh um not sure

OpenStudy (austinl):

Think of it like this, \(0.02x+70<0.09x\Rightarrow~0.02x+70=0.09x\) What would you do first to solve for that?

OpenStudy (anonymous):

um add 70 to both sides?

OpenStudy (austinl):

It is already an addition, so what do you do to cancel out an addition?

OpenStudy (anonymous):

subtract

OpenStudy (austinl):

Well, rewind, we want terms with variables on one side and constants on the other. We can do this in one step without messing with the 70. But on the same track, we would subtract 0.02x from both sides! Which when done will look like this, \(70<0.07x\) Now what do we do to cancel out a multiplication?

OpenStudy (anonymous):

um divide

OpenStudy (austinl):

Correctomundo, so therefore, \(\displaystyle x>\frac{70}{0.07}\) Which is what?

OpenStudy (anonymous):

1000

OpenStudy (austinl):

So that means that she would have to copy 1000 sheets to make copier 2 more cost effective.

OpenStudy (anonymous):

oh ok thats it

OpenStudy (austinl):

Woot woot! Good job!

OpenStudy (anonymous):

thx